OpenStudy (anonymous):

Sharmi ran 3 miles in 21 minutes. How long would it take her to run 5 miles? 12 minutes 27 minutes 35 minutes 42 minutes

OpenStudy (anonymous):

do the same thing, that u did in the last problem, 3*7=21. What is 5*7?

OpenStudy (zpupster):

set up ratio 3/21 and 5/? cross multiply. 3?=5(21) 105/3 35

OpenStudy (anonymous):

either way u do it, the answer is 35.

OpenStudy (zpupster):

you bet